# Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging

### v2

@@ Abstract
2-boson interactions as caging mechanisms.— Interestingly, our density interaction $U$ can also act as a non-local cage when it is situated away from the boundary site $x=1$. In Fig. 3, we demonstrate how this interaction largely confines the evolved state to the left of $x_{0}$ i.e. $x\leq x_{0}$, if the initial state has both bosons in the same region. And, more elaborately, for an initial state with both bosons lying to the right of $x_{0}$, Fig. 4 shows how the interaction will first act as a barrier to the evolved state, but then eventually still trap the particle flux that leaked through it due to NHSE pumping. Overall, this caging mechanism hence further enhances the ultra-strong boson condensation at the boundary. We shall fix the density interaction to be at $x_{0}=3$, and consider initial states $(2,2)$ and $(4,4)$ in Fig. 3 and Fig. 4 respectively.
@@
This caging mechanism is most intuitively represented in the 2D configuration space $(x,x^{\prime})$, where $x$ and $x^{\prime}$ represent the positions of the two bosons [^1] Here $(x,x^{\prime})$ and $(x^{\prime},x)$ refer to the same state due to bosonic statistics [72, 132, 133, 134, 75, 135].. A key observation is that the nonlinear term $\frac{1}{2}Un_{x_{0}}^{2}$ can be interpreted as a non-local L-shaped “cage” [Figs. 3(a) and 4(a)], with potential “walls” [72, 75] of height $\frac{1}{2}U$ across the entire lines $(x,x_{0})$ and $(x_{0},x)$ where only one boson is at $x_{0}$. These walls cross at $(x_{0},x_{0})$, where double bosonic occupancy at $x_{0}$ gives an even higher potential barrier of $2U$.
@@
Fig. 3(a) compares the dynamical evolution of an initial state $(2,2)$ in the (a1) absence and (a2) presence of the density interaction $U$. In Fig. 3(a1) with $U=0$, the smoothed probability cloud $\bar{P}_{xx^{\prime}}(t)$ [Eq. 6] of finding the bosons at $(x,x^{\prime})$ spreads out freely and reflects against the boundaries, even though it is slightly amplified towards the $(1,1)$ corner due to leftwards NHSE pumping. However, in Fig. 3(a2), this spreading is drastically contained within the $3\times 3$ L-shaped cage formed by nonzero $U$ at $x_{0}=3$ [see Supplemental videos]. This caging, which requires both the $U$ interaction and directed amplification towards $(1,1)$, indeed leads to far enhanced boundary boson condensation $\bar{P}_{11}(t)$, as shown in the red curve in Fig. 3(b1).
@@
Interestingly, the caging mechanism protects the $\bar{P}_{11}$ condensation even when the initial state is outside the cage. In Fig. 4(a) with both bosons initially to the right of $x_{0}=3$ at $(4,4)$, their flux $\bar{P}_{xx^{\prime}}(t)$ still gradually enters the $3\times 3$ cage due to the NHSE pumping towards $(1,1)$. Comparing the non-interacting (a1) with the interacting (a2) cases between $t=0$ to $2.2$, the interaction-induced potential walls only slow the diffusion into the cage slightly. However, once the bosons have entered the cage, they are subject to the same trapping mechanism as in Fig. 3, thereby also experiencing eventual ultra-strong condensation at $x=1$ [red in Fig. 4(b1)].
@@
Notably, this caging mechanism is an emergent consequence of few-body density interactions, and cannot be replicated by an effective potential barrier in the single-particle context. Plotted in Figs. 3(b2) and 4(b2) are the evolutions of the densities $\bar{\rho}_{x=1}$ of a single boson in the same 1D SSH chain of Eq. 1, but with an on-site potential $U$ at $x_{0}=3$. Evidently, the $U\neq 0$ cases (orange) do not trap the boson at $x=1$ boundary any more than the $U=0$ cases (blue), with the boundary boson density remaining far lower than the $\bar{P}_{11}(t)$ of the interacting NHSE 2-boson cases [red in Figs. 3(b1) and 4(b1)].
@@
Generalizations.— Although we have explicitly considered only two particles, the caging mechanism that leads to ultra-strong boundary localization holds for rather general density interactions with arbitrary numbers of bosons. With $N$ bosons, each Fock state is indexed by a lattice point $\vec{x}=(x_{1},...,x_{N})^{T}$ living in the configuration space $\mathbb{R}^{N}$ [84, 85, 86, 87], and the density takes distinct integer values in the various hyperplanes, depending on the bosonic occupancy. Given a generic interaction that is a multinomial in the densities $n_{1},n_{2},...$ at various sites, a monomial $Un_{x_{1}}n_{x_{2}}...n_{x_{M}}$ gives different energy offset in the corresponding hyperplanes, such as to form the barriers of a “hyper-cage”. For instance, with 4 bosons, an interaction $Un_{x_{0}}n_{x^{\prime}_{0}}$ gives an energy $U$ in the planes where two bosons at fixed at $x_{0}$ and $x^{\prime}_{0}$, $2U$ along lines where three bosons are fixed at either $x_{0}$ or $x^{\prime}_{0}$, etc. Due to the non-locality of the NHSE [22, 23, 14, 24, 25], caging can occur as long as translation invariance is broken in the configuration space, even if the interaction-induced barriers do not completely enclose any region.
Discussion.— We have revealed how the interplay between the NHSE, topology and boson-boson interactions can unexpectedly lead to ultra-strong bosonic condensation. This particle condensation strongly exceeds that of NHSE-pumped free bosons, being also facilitated by the interaction-induced hybridization of topological and bulk states, as well as an emergent non-local caging mechanism. Our findings generalize to higher numbers of bosons as well as generic density interactions and topologies, bringing forth a new approach to trapping and controlling bosons. Experimentally, demonstrations of non-Hermitian systems have been maturely built upon mechanical arrays [88, 89], photonic [90, 91, 62, 92, 29], electrical circuit [24, 55, 93, 81] and quantum circuit platforms [7]. Various platforms can potentially implement the setup for locally interacting bosons to incorporate the $n_{x}^{2}$ term, including optical lattices [94, 95, 96, 97, 98, 99], photonic systems [100, 101, 102, 103, 104, 105, 106, 107, 73], circuit QED [108, 109, 110, 111, 112],various quantum simulators [113, 114, 115, 71, 116, 117, 97, 118, 119, 120, 121, 122, 123, 124, 109, 125, 126, 127, 128, 129, 130, 131] as well as photonic resonator arrays, as further elaborated in [78].
Supplemental Material for “Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging”
Mengjie Yang
Luqi Yuan
Ching Hua Lee
October 14, 2024
S1 Single-boson and two-boson spectra
@@
The single-boson band structure, shown in Fig. S1(a), contains two symmetrically gapped bulk bands ($E_{1}$ and $E_{2}$) separated by in-gap topological zero modes ($E_{0}$) at zero energy. The corresponding eigenstates demonstrate distinct localization profiles, with boundary skin-localized states for bulk bands [Fig. S1(b)] as well as topological zero modes [Fig. S1(c)].
@@
Figure S1: Single-boson spectrum of the non-Hermitian SSH model [Eq. S1 with $U=0$] under open boundary conditions with $t_{L}=1.6$, $t_{R}=0.4$ and $t_{0}=3$. (a) The energy spectrum shows two bulk bands ($E_{1}$ and $E_{2}$) separated by topological zero modes ($E_{0}$). (b) Localization profile $|\psi|^{2}$ of left boundary-localized states in the bulk bands. (c) Localization of the topological zero mode, also confined at the left boundary.
For the two-boson case in the non-interacting limit, the energy spectrum can be systematically constructed through all possible pairwise combinations of the single-particle energies. As shown in Fig. S2(a), this results in five distinct bands rather than three. The lowest [band 1 in Fig. S2] and highest [band 5 in Fig. S2] bands correspond to configurations where both bosons occupy the lower ($E_{1}\oplus E_{1}$) or upper ($E_{2}\oplus E_{2}$) bulk bands, respectively. The intermediate bands 2 and 4 in Fig. S2 arise from mixed occupations, where one boson occupies the zero mode while the other occupies either the lower ($E_{0}\oplus E_{1}$) or upper ($E_{0}\oplus E_{2}$) bulk band.
@@
Figure S2: Two-boson spectrum in the non-interacting limit ($U=0$). (a) Energy spectrum showing five bands arising from combinations of single-particle energies. Band 1 results from hybridization of $E_{1}\oplus E_{1}$. Band 2 results from hybridization of $E_{1}(E_{0})\oplus E_{0}(E_{1})$. Band 3 results from hybridization of $E_{1}\oplus E_{2}$ and $E_{0}\oplus E_{0}$ configurations. Band 4 results from hybridization of $E_{2}(E_{0})\oplus E_{0}(E_{2})$. Band 5 results from hybridization of $E_{2}\oplus E_{2}$. (b) Localization pattern of a zero-corner state in band 3. (c) Localization of mixed bulk-topological states in bands 2 and 4, where either $x$ or $x^{\prime}$ (both not both) are boundary-localized.
@@
Notably, the central band [band 3 in Fig. S2] emerges from the merging of two distinct contributions: one where the bosons are distributed between the lower and upper bulk bands ($E_{1}\oplus E_{2}$), and another (much rarer) instance where both bosons occupy the zero mode ($E_{0}\oplus E_{0}$). This explains why we observe five bands rather than six possible energy band combinations, as the energetic proximity of these configurations leads to their mixing into a single band.
@@
The localization profiles shown in Fig. S2(b) and (c) reflect this hybridization, exhibiting characteristics that can be understood as superpositions of the single-particle localization profiles.
S2 Possible physical realizations
@@
In the main text, we have proposed a bosonic interacting model in a non-Hermitian SSH lattice, as shown in Eq.1, also reproduced here as Eq. S1:
$H=$ $\sum_{x=1}^{L}t_{L}b_{2x-1}^{\dagger}b_{2x}+t_{R}b_{2x}^{\dagger}b_{2x-1}$ (S1)
@@
Below, we discuss how the bosonic interacting model of Eq. S1 may be potentially realized using photonic resonator arrays. Arrays of micro-resonators have been extensively utilized to implement one-dimensional (1D) and two-dimensional (2D) photonic lattices, particularly in the investigation of non-Hermitian photonics over the past decade [136, 137, 138, 139, 140, 141]. These designs predominantly employ ring resonators, which can achieve exceptionally high Q factors [141].
@@
In Ref. [142], the authors proposed how a non-Hermitian SSH model can be realized using photonic coupled resonant arrays. The non-Hermitian asymmetric coupling can be realized through judicious incorporation of optical gain and loss elements into unidirectional coupling link rings.
@@
To implement the nonlinear density term, we can employ the Kerr effect which introduces third-order nonlinearity $\chi^{(3)}$. This nonlinear optical effect causes the refractive index of a medium to change with the intensity of light. The third-order nonlinearity can be achieved through natural materials [143, 144, 145, 146] (although the $\chi^{(3)}$ value is relatively small) or by employing cavity QED based on electromagnetically induced transparency (EIT) to achieve an equivalent $\chi^{(3)}$. The latter involves introducing atoms [77, 147, 148, 149] or quantum dots into a cavity [34, 150, 151, 152, 153] and, through appropriate external field control, inducing a target optical field with strong effective nonlinearity.
S3 Bosonic condensation in the Hatano-Nelson model
In the main text, we have discussed the emergence of ultra-strong bosonic condensation in the non-Hermitian SSH model. Here, we show that the ultra-condensation phenomenon and the caging mechanism are not exclusive to the non-Hermitian SSH model but can also be observed in the Hatano-Nelson model. The Hatano-Nelson model is known to exhibit a non-Hermitian skin effect, where the bulk states are localized at the edges of the system [154]. After including the bosonic interactions, the interacting Hatano-Nelson Hamiltonian is given by
$H_{\text{HN}}=\sum_{x=1}^{L-1}t_{L}b_{x}^{\dagger}b_{x+1}+t_{R}b_{x+1}^{\dagger}b_{x}+\frac{U}{2}n_{x_{0}}^{2},$ (S2)
where $t_{L}$ and $t_{R}$ are the left and right hopping amplitudes, respectively. And $U$ is the on-site interaction strength. We will use the same measurement as in the main text—specifically, the smoothed two-boson correlation probability, $\bar{P}_{xx^{\prime}}(t)$, from Eq. 6 in the main text, which represents the trends of the probability of observing bosons at sites $x$ and $x^{\prime}$ simultaneously at time $t$—to demonstrate ultra-strong bosonic condensation in the Hatano-Nelson model.
In Fig. S3, we show the smoothed two-boson correlation probability $\bar{P}_{11}(t)$ and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the Hatano-Nelson model. We consider two different initial states similar as in the main text: one where both bosons are initially inside the cage $(1,1)$ and another where both bosons are initially outside the cage $(3,3)$. In both cases, we found the predominantly red curves (non-Hermitian, Interacting) exhibit the same ultra-strong bosonic condensation and caging behavior as in the Hatano-Nelson model. This demonstrates that the ultra-condensation phenomenon and the caging mechanism are not exclusive to the non-Hermitian SSH model but can be observed in other non-Hermitian models as well.
@@
Figure S3: (a) and (b) show the smoothed two-boson correlation probability $P_{11}(t)$ in Eq. 6 in the main text and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the initial state $(1,1)$ [inside the cage] at interaction position $x_{0}=2$. (c) and (d) show the same quantities for the initial state $(3,3)$ [outside the cage] and $x_{0}=2$. $r=t_{L}/t_{R}=4,t_{L}=1/t_{R}=2,U=6,L=10$ for the Hatano-Nelson model in Eq. S2.
S4 Bosonic condensation in the non-Hermitian SSH model with trivial topology
In the main text, we discussed the emergence of ultra-strong bosonic condensation in the non-Hermitian SSH model with nontrivial topology, where $t_{L}\times t_{R}=0.4\times 1.6<t_{0}^{2}=3^{2}$. Here, we demonstrate that both the ultra-condensation phenomenon and the caging mechanism can also be observed in the non-Hermitian SSH model with trivial topology, where $t^{\prime}_{L}\times t^{\prime}_{R}>t_{0}^{\prime 2}$.
For this section, we use the parameters $t_{0}^{\prime}=1$, $t^{\prime}_{L}=4.8$, and $t^{\prime}_{R}=1.2$. These values are chosen such that $t_{0}(t_{L}+t_{R})=t_{0}^{\prime}(t^{\prime}_{L}+t^{\prime}_{R})$ and $r=t^{\prime}_{L}/t^{\prime}_{R}=4$, ensuring similar scales and hopping asymmetry for both the nontrivial and trivial cases. Here, $t^{\prime}_{L}\times t^{\prime}_{R}=4.8\times 1.2>t_{0}^{\prime 2}=1$, corresponding to the trivial case. The same measurement as in the main text—specifically, the smoothed two-boson correlation probability, $\bar{P}_{xx^{\prime}}(t)$, from Eq. 6 in the main text—is used to demonstrate ultra-strong bosonic condensation in the non-Hermitian SSH model with trivial topology. And $\bar{P}_{\text{cage}}(t)$, the smoothed two-boson probability within the $x_{0}\times x_{0}$ cage, is used to illustrate the caging mechanism.
@@
Figure S4: (a) and (b) show the smoothed two-boson correlation probability $P_{11}(t)$ in Eq. 6 in the main text and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the initial state $(1,1)$ [inside the cage] at interaction position $x_{0}=2$. (c) and (d) show the same quantities for the initial state $(3,3)$ [outside the cage] and $x_{0}=2$. $r=t_{L}/t_{R}=4,t_{L}=1/t_{R}=2,U=6,L=10$ for the non-Hermitian SSH model in Eq. 1 in the main text with $t_{0}=t_{0}^{\prime}=1$, $t_{L}=t^{\prime}_{L}=4.8$, and $t_{R}=t^{\prime}_{R}=1.2$.
The ultra-strong bosonic condensation and caging behavior are illustrated in Fig. S4. We also consider two different initial states: one where both bosons are initially inside the cage $(1,1)$ and another where both bosons are initially outside the cage $(3,3)$. In both scenarios, we found that the predominantly red curves (non-Hermitian, interacting) exhibit the same ultra-strong bosonic condensation and caging behavior as observed in the non-Hermitian SSH model with nontrivial topology. This demonstrates that the ultra-condensation phenomenon and the caging mechanism are not exclusive to the non-Hermitian SSH model with nontrivial topology; rather, they can also be observed in the non-Hermitian SSH model with trivial topology.
The key difference between the trivial and nontrivial models, as discussed in the main text, lies in the absence of a topological origin in the trivial case. In this case, the topological bands, specifically bands 2 and 4, are no longer present, as illustrated in Fig. S5(b). Instead, the bulk states are localized at the edges, which is also shown in Fig. S5(b). The ultra-strong condensation described in the main text, which has a topological origin in the nontrivial case, is absent in the trivial case. Furthermore, the ultra-strong condensation shown in Fig. S5(a) does not correspond to the boundary localization of the 2-boson spectral bands in the trivial model.

### v3

@@ III.3 Two-boson interactions as caging mechanisms
Interestingly, our density interaction $U$ can also act as a non-local cage when it is situated away from the boundary site $x=1$. In Fig. 3, we demonstrate how this interaction largely confines the evolved state to the left of $x_{0}$ i.e. $x\leq x_{0}$, if the initial state has both bosons in the same region. And, more elaborately, for an initial state with both bosons lying to the right of $x_{0}$, Fig. 4 shows how the interaction will first act as a barrier to the evolved state, but then eventually still trap the particle flux that leaked through it due to NHSE pumping. Overall, this caging mechanism hence further enhances the ultra-strong boson clustering at the boundary. We shall fix the density interaction to be at $x_{0}=3$, and consider initial states $(2,2)$ and $(4,4)$ in Fig. 3 and Fig. 4 respectively.
@@
This caging mechanism is most intuitively represented in the 2D configuration space $(x,x^{\prime})$, where $x$ and $x^{\prime}$ represent the positions of the two bosons [^1] Here $(x,x^{\prime})$ and $(x^{\prime},x)$ refer to the same state due to bosonic statistics [33, 41, 42, 43, 37].. A key observation is that the nonlinear term $\frac{1}{2}Un_{x_{0}}^{2}$ can be interpreted as a non-local L-shaped “cage” [Figs. 3(a) and 4(a)], with potential “walls” [33, 41, 42, 43, 37] of height $\frac{1}{2}U$ across the entire lines $(x,x_{0})$ and $(x_{0},x)$ where only one boson is at $x_{0}$. These walls cross at $(x_{0},x_{0})$, where double bosonic occupancy at $x_{0}$ gives an even higher potential barrier of $2U$.
@@
Fig. 3(a) compares the dynamical evolution of an initial state $(2,2)$ in the (a1) absence and (a2) presence of the density interaction $U$. In Fig. 3(a1) with $U=0$, the smoothed probability cloud $\bar{P}_{xx^{\prime}}(t)$ [Eq. 6] of finding the bosons at $(x,x^{\prime})$ spreads out freely and reflects against the boundaries, even though it is slightly amplified towards the $(1,1)$ corner due to leftwards NHSE pumping. However, in Fig. 3(a2), this spreading is drastically contained within the $3\times 3$ L-shaped cage formed by nonzero $U$ at $x_{0}=3$ [see Supplemental videos]. This caging, which requires both the $U$ interaction and directed amplification towards $(1,1)$, indeed leads to far enhanced boundary boson clustering $\bar{P}_{11}(t)$, as shown in the red curve in Fig. 3(b1).
@@
Interestingly, the caging mechanism protects the $\bar{P}_{11}$ clustering even when the initial state is outside the cage. In Fig. 4(a) with both bosons initially to the right of $x_{0}=3$ at $(4,4)$, their flux $\bar{P}_{xx^{\prime}}(t)$ still gradually enters the $3\times 3$ cage due to the NHSE pumping towards $(1,1)$. Comparing the non-interacting (a1) with the interacting (a2) cases between $t=0$ to $2.2$, the interaction-induced potential walls only slow the diffusion into the cage slightly. However, once the bosons have entered the cage, they are subject to the same trapping mechanism as in Fig. 3, thereby also experiencing eventual ultra-strong clustering at $x=1$ [red in Fig. 4(b1)].
@@
Notably, this caging mechanism is an emergent consequence of few-body density interactions, and cannot be replicated by an effective potential barrier in the single-particle context. Plotted in Figs. 3(b2) and 4(b2) are the evolutions of the densities $\bar{\rho}_{x=1}$ of a single boson in the same 1D SSH chain of Eq. 1, but with an on-site potential $U$ at $x_{0}=3$. The $U\neq 0$ cases (orange) do not trap the boson at $x=1$ boundary any more than the $U=0$ cases (blue), with the boundary boson density remaining far lower than the $\bar{P}_{11}(t)$ of the interacting NHSE 2-boson cases [red in Figs. 3(b1) and 4(b1)].
@@
IV Discussion and generalizations
Although we have explicitly considered only two particles, the caging mechanism that leads to ultra-strong boundary clustering holds for rather general density interactions with arbitrary numbers of bosons. With $N$ bosons, each Fock state is indexed by a lattice point $\vec{x}=(x_{1},...,x_{N})^{T}$ living in the configuration space $\mathbb{R}^{N}$ [44, 45, 46, 47], and the density takes distinct integer values in the various hyperplanes, depending on the bosonic occupancy. Given a generic interaction that is a multinomial in the densities $n_{1},n_{2},...$ at various sites, a monomial $Un_{x_{1}}n_{x_{2}}...n_{x_{M}}$ gives different energy offset in the corresponding hyperplanes, such as to form the barriers of a “hyper-cage”. For instance, with 4 bosons, an interaction $Un_{x_{0}}n_{x^{\prime}_{0}}$ gives an energy $U$ in the planes where two bosons at fixed at $x_{0}$ and $x^{\prime}_{0}$, $2U$ along lines where three bosons are fixed at either $x_{0}$ or $x^{\prime}_{0}$, etc. Due to the non-locality of the NHSE [2, 11, 48, 49], caging can occur as long as translation invariance is broken in the configuration space, even if the interaction-induced barriers do not completely enclose any region.
We have revealed how the interplay between the NHSE, topology, and boson-boson interactions can unexpectedly lead to ultra-strong bosonic clustering. This particle clustering strongly exceeds that of NHSE-pumped free bosons, being also facilitated by the interaction-induced hybridization of topological and bulk states and an emergent non-local caging mechanism. Our findings generalize to higher numbers of bosons as well as generic density interactions and topologies, bringing forth a new approach to trapping and controlling bosons. Experimentally, demonstrations of non-Hermitian systems have been maturely built upon mechanical arrays [50, 51], photonic [52, 53], electrical circuit [48, 39] and quantum circuit platforms [6]. Various platforms can potentially implement the setup for locally interacting bosons to incorporate the $n_{x}^{2}$ term, including optical lattices [54, 55], photonic systems [56, 35], circuit QED [57, 58], various quantum simulators [59, 60, 61, 62] as well as photonic resonator arrays, as further elaborated in Appendix.
Appendix A Single-boson and two-boson spectra
@@
The single-boson band structure, shown in Fig. 5(a), contains two symmetrically gapped bulk bands ($E_{1}$ and $E_{2}$) separated by in-gap topological zero modes ($E_{0}$) at zero energy. The corresponding eigenstates demonstrate distinct localization profiles, with boundary skin-localized states for bulk bands [Fig. 5(b)] as well as topological zero modes [Fig. 5(c)].
@@
Figure 5: Single-boson spectrum of the non-Hermitian SSH model [Eq. 7 with $U=0$] under open boundary conditions with $t_{L}=1.6$, $t_{R}=0.4$ and $t_{0}=3$. (a) The energy spectrum shows two bulk bands ($E_{1}$ and $E_{2}$) separated by topological zero modes ($E_{0}$). (b) Localization profile $|\psi|^{2}$ of left boundary-localized states in the bulk bands. (c) Localization of the topological zero mode, is also confined at the left boundary.
@@
Figure 6: Two-boson spectrum in the non-interacting limit ($U=0$). (a) Energy spectrum showing five bands arising from combinations of single-particle energies. Band 1 results from hybridization of $E_{1}\oplus E_{1}$. Band 2 results from hybridization of $E_{1}(E_{0})\oplus E_{0}(E_{1})$. Band 3 results from the hybridization of $E_{1}\oplus E_{2}$ and $E_{0}\oplus E_{0}$ configurations. Band 4 results from hybridization of $E_{2}(E_{0})\oplus E_{0}(E_{2})$. Band 5 results from hybridization of $E_{2}\oplus E_{2}$. (b) Localization pattern of a zero-corner state in band 3. (c) Localization of mixed bulk-topological states in bands 2 and 4, where either $x$ or $x^{\prime}$ (both not both) are boundary-localized.
For the two-boson case in the non-interacting limit, the energy spectrum can be systematically constructed through all possible pairwise combinations of the single-particle energies. As shown in Fig. 6(a), this results in five distinct bands rather than three. The lowest [band 1 in Fig. 6] and highest [band 5 in Fig. 6] bands correspond to configurations where both bosons occupy the lower ($E_{1}\oplus E_{1}$) or upper ($E_{2}\oplus E_{2}$) bulk bands, respectively. The intermediate bands 2 and 4 in Fig. 6 arise from mixed occupations, where one boson occupies the zero mode while the other occupies either the lower ($E_{0}\oplus E_{1}$) or upper ($E_{0}\oplus E_{2}$) bulk band.
@@
Notably, the central band [band 3 in Fig. 6] emerges from the merging of two distinct contributions: one where the bosons are distributed between the lower and upper bulk bands ($E_{1}\oplus E_{2}$), and another (much rarer) instance where both bosons occupy the zero mode ($E_{0}\oplus E_{0}$). This explains why we observe five bands rather than six possible energy band combinations, as the energetic proximity of these configurations leads to their mixing into a single band.
@@
The localization profiles shown in Fig. 6(b) and (c) reflect this hybridization, exhibiting characteristics that can be understood as superpositions of the single-particle localization profiles.
Appendix B Possible physical realizations
@@
In the main text, we have proposed a bosonic interacting model in a non-Hermitian SSH lattice, as shown in Eq.1, also reproduced here as Eq. 7:
$H=$ $\sum_{x=1}^{L}t_{L}b_{2x-1}^{\dagger}b_{2x}+t_{R}b_{2x}^{\dagger}b_{2x-1}$ (7)
@@
Below, we discuss how the bosonic interacting model of Eq. 7 may be potentially realized using photonic resonator arrays. Arrays of micro-resonators have been extensively utilized to implement one-dimensional (1D) and two-dimensional (2D) photonic lattices, particularly in the investigation of non-Hermitian photonics over the past decade [63, 64, 65]. These designs predominantly employ ring resonators, which can achieve exceptionally high Q factors [64].
@@
In Ref. [66], the authors proposed how a non-Hermitian SSH model can be realized using photonic coupled resonant arrays. The non-Hermitian asymmetric coupling can be realized through the judicious incorporation of optical gain and loss elements into unidirectional coupling link rings.
@@
To implement the nonlinear density term, we can employ the Kerr effect which introduces third-order nonlinearity $\chi^{(3)}$. This nonlinear optical effect causes the refractive index of a medium to change with the intensity of light. The third-order nonlinearity can be achieved through natural materials [67, 68] (although the $\chi^{(3)}$ value is relatively small) or by employing cavity QED based on electromagnetically induced transparency (EIT) to achieve an equivalent $\chi^{(3)}$. The latter involves introducing atoms [34, 69] or quantum dots into a cavity [17, 70, 71, 72] and, through appropriate external field control, inducing a target optical field with strong effective nonlinearity.
Appendix C Bosonic clustering in the Hatano-Nelson model
@@
Figure 7: (a) and (b) show the smoothed two-boson correlation probability $P_{11}(t)$ in Eq. 6 in the main text and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the initial state $(1,1)$ [inside the cage] at interaction position $x_{0}=2$. (c) and (d) show the same quantities for the initial state $(3,3)$ [outside the cage] and $x_{0}=2$. $r=t_{L}/t_{R}=4,t_{L}=1/t_{R}=2,U=6,L=10$ for the Hatano-Nelson model in Eq. 8.
@@
Figure 8: (a) and (b) show the smoothed two-boson correlation probability $P_{11}(t)$ in Eq. 6 in the main text and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the initial state $(1,1)$ [inside the cage] at interaction position $x_{0}=2$. (c) and (d) show the same quantities for the initial state $(3,3)$ [outside the cage] and $x_{0}=2$. $r=t_{L}/t_{R}=4,t_{L}=1/t_{R}=2,U=6,L=10$ for the non-Hermitian SSH model in Eq. 1 in the main text with $t_{0}=t_{0}^{\prime}=1$, $t_{L}=t^{\prime}_{L}=4.8$, and $t_{R}=t^{\prime}_{R}=1.2$.